the square of a number is 255greater than twice the number. what is the number?
Answer by faceoff57(108)

the square of a number is 255greater than twice the number. what is the number?
X^2 = 255 + 2x
X^2 - 2x - 255 = 0
(X + 15)(x - 17) = 0
X + 15 = 0
X = -15
X - 17 = 0
X = 17
-15 or 17